OpenStudy (anonymous):

how do you factor 5^3 +5c

OpenStudy (unklerhaukus):

\[5^3+5c=5(5^2+c) \]

OpenStudy (anonymous):

\[5^3+5c\] would not help i think my frnd u missed out some thing!

OpenStudy (anonymous):

It should of been 5c^3 +5c

OpenStudy (unklerhaukus):

\[5c^3 +5c=5c(c^2+1)\]